# When Personas Override Payoffs: Role Identity Bias in Multi-Agent LLM Decision-Making

### v1

@@ Appendix A Experimental Design
In the persona condition, each of the four agents receives a role-specific description reflecting realistic motivations associated with that stakeholder (e.g., profitability concerns for Industrialist, public welfare tradeoffs for Government). In the no-persona baseline condition, agents receive no role description and are instructed only to select actions based on best response Nash equilibrium reasoning. This manipulation isolates role identity as a causal factor. We also compare two settings where payoffs are either visible or hidden. In the hidden payoff condition, agents do not receive explicit payoff matrices but infer incentives from the scenario description. In the visible payoff condition, agents receive explicit payoff matrices and are instructed to identify a Nash equilibrium directly. This allows us to test whether explicit incentive information mitigates or amplifies role identity bias. All experiments use chain-of-thought reasoning instructions, enabling us to understand how models balance identity-driven behavior and payoff-optimal reasoning.
Appendix B Model Versions and Hyperparameters
B.1. Model Versions
All experiments were conducted using the following model versions:
B.2. Generation Hyperparameters
  • Temperature: 0.2 (low temperature for more deterministic outputs)
  • Top-p (nucleus sampling): 0.9
  • Max new tokens: 256 tokens per generation
  • Sampling: Enabled (do_sample=True) with temperature ¿ 0
B.3. Experimental Repetitions
@@
Appendix C Experimental Prompts and Persona Descriptions
@@
C.1. Action Choice Prompt
@@ C.2. Persona Descriptions
C.3. Neutral Agent Configuration
@@ C.4. Action Mappings
Appendix D Scenario Examples
This section provides example scenarios from each scenario type used in the experiments. Scenarios are classified as “economic” (where Tragedy of the Commons is the Nash equilibrium), “environmental” (where Green Transition is the Nash equilibrium) or Mixed/Complex scenarios (where economic and environmental factors are balanced) based on the payoff structure of the underlying game. Table D shows an example of each type of scenario, and Table 8 shows subtypes of scenarios.
@@ Appendix D Scenario Examples
Appendix E Payoff Matrix Format
@@
E.1. Example Payoff Matrix
@@
Appendix F Chain-of-Thought Keyword Categories
@@
F.1. Example Chain-of-Thought Rationales
@@
Appendix G Supplementary Results
@@
Table 11. Overall Nash Equilibrium Rates: Comparison of Hidden vs Visible Payoffs. This table compares overall Nash equilibrium rates (across all scenarios) between hidden and visible payoffs conditions. Each row shows a different model-persona combination. The table reveals that with personas, hidden payoffs achieve equal or higher Nash rates than visible payoffs, suggesting reasoning scaffolding is more effective when role identity bias is present. Without personas, visible payoffs show large decreases for Mistral and Qwen models, but for Qwen, this is due to enabling Tragedy in economic scenarios (which is correct behavior).
@@
Appendix H Chain-of-Thought Analysis Figures
@@
H.1. Keyword Patterns Across Conditions
@@
Figure 4. Green Transition Action Selection Rates in Economic Scenarios. This grouped bar chart shows the percentage of action profiles where models select Green Transition actions (socially preferred but payoff-suboptimal) in economic scenarios where Tragedy of Commons is the payoff-optimal equilibrium. When personas are present, all models show substantial Green Transition selection rates (20-59%) despite Tragedy being payoff-optimal, demonstrating that role identity bias operates at the action level even when models fail to coordinate on Nash equilibria. Qwen models show the highest Green Transition rates with personas (40-59%), while Llama and Mistral show moderate rates (20-30%). Removing personas substantially reduces Green Transition selection for all models, with rates dropping to near-zero when payoffs are visible and personas are absent, confirming that personas drive the preference for socially aligned actions. The persistence of Green Transition selection with personas, even in conditions where Nash equilibrium rates are zero, indicates that role identity bias affects individual action choices independently of coordination success, revealing a pervasive influence of persona-driven reasoning on strategic decision-making.
Figure 5. CoT Keyword Frequency Patterns by Model and Condition. This figure shows keyword frequency distributions (normalized per 1000 rationales) across all four experimental conditions for each model.
H.2. Reasoning Shift from Identity-Driven to Payoff-Optimal
@@
H.3. Model Comparison of Game-Theoretic Reasoning
@@
H.4. Strategic Keywords: Unique Pattern in Qwen 32B

### v2

@@ Appendix A Experimental Design
In the persona condition, each of the four agents receive a role-specific description reflecting realistic motivations associated with that stakeholder (e.g., profitability concerns for Industrialist, public welfare tradeoffs for Government). In the no-persona baseline condition, agents receive no role description and are instructed only to select actions based on best response Nash equilibrium reasoning. This manipulation isolates role identity as a causal factor. We also compare two settings where payoffs are either visible or hidden. In the hidden payoff condition, agents do not receive explicit payoff matrices but infer incentives from the scenario description. In the visible payoff condition, agents receive explicit payoff matrices and are instructed to identify a Nash equilibrium directly. This allows us to test whether explicit incentive information mitigates or amplifies role identity bias. All experiments use chain-of-thought reasoning instructions, enabling us to understand how models balance identity-driven behavior and payoff-optimal reasoning.
A.1. Experimental Repetitions
@@
A.2. Prompts and Persona Descriptions
@@
A.3. Action Choice Prompt
@@ A.4. Persona Descriptions
A.5. Neutral Agent Configuration
@@ A.6. Action Mappings
A.7. Scenario Examples
This section provides example scenarios from each scenario type used in the experiments. Scenarios are classified as “economic” (where Tragedy of the Commons is the Nash equilibrium), “environmental” (where Green Transition is the Nash equilibrium) or Mixed/Complex scenarios (where economic and environmental factors are balanced) based on the payoff structure of the underlying game. Table A.7 shows an example of each type of scenario, and Table 8 shows subtypes of scenarios.
@@ A.7. Scenario Examples
A.8. Payoff Matrix Format
@@
A.8.1. Example Payoff Matrix
@@
Appendix B Model Versions and Hyperparameters
B.1. Model Versions
All experiments were conducted using the following model versions:
B.2. Generation Hyperparameters
  • Temperature: 0.2 (low temperature for more deterministic outputs)
  • Top-p (nucleus sampling): 0.9
  • Max new tokens: 256 tokens per generation
  • Sampling: Enabled (do_sample=True) with temperature ¿ 0
Appendix C Chain-of-Thought Keyword Categories
@@
C.1. Example Chain-of-Thought Rationales
@@
Appendix D Supplementary Results
@@
Table 11. Overall Nash Equilibrium Rates: Comparison of Hidden vs Visible Payoffs. This table compares overall Nash equilibrium rates (across all scenarios and models) between hidden and visible payoffs conditions. Each row shows a different model-persona combination. The table reveals that with personas, hidden payoffs achieve equal or higher Nash rates than visible payoffs, suggesting reasoning scaffolding is more effective when role identity bias is present. Without personas, visible payoffs show large decreases for Mistral and Qwen models, but for Qwen, this is due to enabling Tragedy in economic scenarios (which is correct behavior).
@@
Appendix E Chain-of-Thought Analysis Figures
@@
E.1. Keyword Patterns Across Conditions
@@
Figure 5. CoT Keyword Frequency Patterns by Model and Condition. This figure shows keyword frequency distributions (normalized per 1000 rationales) across all four experimental conditions for each model, detailing the distribution of keywords within the generated rationales.
E.2. Reasoning Shift from Identity-Driven to Payoff-Optimal
@@
E.3. Model Comparison of Game-Theoretic Reasoning
@@
E.4. Strategic Keywords - A Unique Pattern in Qwen 32B